OpenStudy (lxelle):

A trader prepares his accounts annually to 30 April. He pays annual rent of $12000 and makes the payments quarterly in advance on 1 Jan, 1 Apr, 1 July and 1 October. Which amount should be included in his accounts for the year ended 30 April 2011? A. $1000 Accrual B. $1000 Prepayment C. $2000 Accrual D. $2000 Prepayment

OpenStudy (anonymous):

=12000/4 = 3000 this is the amount of his quarterly payments year is May 2010 to Aprll 2011 = 12000/12 = 1000 monthly payment Question states payments are made in Jan Apr July and Oct When that April 2011 payment is made - that reflects 2 months of prepaid

OpenStudy (anwaarullah):

answer is A